# On circumcentered direct methods for monotone variational inequality problems

### v1

@@ 5.1 Examples of operators in ${\rm VIP}(F,C)$
Example 3: Finally, we consider cases with monotone operators which are not paramonotone. $F$ is still defined by (5.9), and $A_{1}\in\mathds{R}^{n_{1}\times n_{1}}$ is also a symmetric positive semidefinite matrix, but $A_{2}\in\mathds{R}^{n_{2}\times n_{2}}$ in (5.10) is skew-symmetric, which means that
$A_{2}^{t}=-A_{2}$ (5.11)
Hence, ${\rm rank}(A+A^{t})<{\rm rank}(A)$. Note that $A$ is the Jacobian matrix of $F$. It is known that if $J(x)$ is the Jacobian matrix of a paramonotone differentiable operator, then ${\rm rank}(J+J^{t})={\rm rank}(J)$, so that the operator $F$ is not paramonotone in this case. Observe that $F$ is monotone, since operators with skew-symmetric Jacobian matrices are trivially monotone.
For each example, we list our results in two tables corresponding to both kinds of comparisons. Finally, we summary the result in all experiments in Figure 1.
Table 1: Median of CPU time and iteration number of Example 1
$n$ $m$ Alg1 Alg2 Alg3 Alg4
CPU(s) Iter CPU(s) Iter CPU(s) Iter CPU(s) Iter
5 2 0.031 908.5 0.141 1670 0.046 1045.5 0.156 1670.5
5 5 0.062 787 1.328 8499.5 1.164 15507 2.921 4320
5 10 0.132 860 8.414 30001 4.376 30001 14.078 10179
10 2 0.031 985 0.289 2991 0.062 1223 0.304 2991.5
10 5 0.063 880 3.250 19006 1.351 16278 1.734 4720
10 10 0.203 1336 7.257 21630 4.421 30001 44.445 26414
20 2 0.070 1519 0.304 2686.5 0.070 1442.5 0.320 2686.5
20 5 0.258 2995 5.085 27223 2.593 30001 20.898 30001
20 10 0.195 1190 3.398 9108.5 2.914 19215 43.087 26634
Table 2: Median of CPU time and error for Example 1
$n$ $m$ Alg1 Alg2 Extra
CPU (s) Error CPU (s) Error CPU (s) Error
5 2 0.031 0.01196 0.539 0.01195 0.523 0.00735
5 5 0.062 0.01026 1.500 0.01003 4.070 0.00724
5 10 0.125 0.00847 8.992 0.00802 23.796 0.00558
10 2 0.031 0.12329 0.187 0.01726 0.671 0.01127
10 5 0.156 0.01239 2.968 0.01244 29.843 0.00804
10 10 0.382 0.01239 9.031 0.01073 82.789 0.00788
20 2 0.054 0.01557 0.398 0.01507 3.085 0.01017
20 5 0.226 0.03118 1.835 0.03116 30.781 0.01906
20 10 0.421 0.03126 9.375 0.02990 171.515 0.02163
Table 3: Median of CPU time and iterations for Example 2
$n$ $m$ Alg1 Alg2 Alg3 Alg4
CPU(s) Iter CPU(s) Iter CPU(s) Iter CPU(s) Iter
5 2 0.031 738.5 0.195 2139.5 0.031 739 0.211 2139.5
5 5 0.062 857 1.906 10824 0.585 6818 7.859 9482
5 10 0.132 878.5 2.406 7737.5 4.429 30001 22.164 12450
10 2 0.031 782.5 0.382 93745 0.032 798.5 1.921 16543
10 5 0.101 1064 1.304 6597 0.250 2367.5 29.218 30001
10 10 0.187 1073 11.101 30001 5.492 30001 67.445 30001
20 2 0.109 3047 0.453 4502 0.125 3146.5 0.546 5093.5
20 5 0.445 2101 4.195 13305 1.945 18460 6.641 13415
20 10 0.437 2673 5.007 7739 53.609 24017 14.562 11119
Table 4: Median of CPU time and error for Example 2
$n$ $m$ Alg1 Alg2 Extra
CPU (s) Error CPU (s) Error CPU (s) Error
5 2 0.031 0.01298 0.321 0.01254 0.617 0.00804
5 5 0.109 0.00982 3.179 0.00962 11.226 0.00602
5 10 0.226 0.01178 9.242 0.01130 44.585 0.01091
10 2 0.148 0.01301 1.273 0.01307 4.968 0.00882
10 5 0.218 0.01431 3.882 0.01431 38.071 0.00932
10 10 0.414 0.01245 9.304 0.01240 122.117 0.00772
20 2 0.093 0.01592 0.375 0.01606 4.821 0.01106
20 5 0.375 0.02437 2.242 0.01948 88.812 0.01356
20 10 0.321 0.01495 8.234 0.01490 125.039 0.01082
Table 5: Median of CPU time and iterations for Example 3
$n$ $m$ Alg1 Alg2 Alg3 Alg4
CPU(s) Iter CPU(s) Iter CPU(s) Iter CPU(s) Iter
5 2 0.031 843.5 0.234 2648.5 0.039 843.5 0.250 2647.5
5 5 0.071 909 1.641 9876.5 2.046 25738 7.734 154749
5 10 0.156 1036 8.578 30001 4.289 30001 50.906 30001
10 2 0.046 1216 0.289 3199.5 0.054 1218 0.304 3199.5
10 5 0.062 762.5 1.195 7284.5 1.210 15399 4.109 5483.5
10 10 0.321 2189.5 8.601 30001 4.289 30001 50.906 30001
20 2 0.046 1394 0.273 2988 0.062 1434 0.281 2998
20 5 0.289 3672.5 1.632 8685 1.757 22185 18.203 21868
20 10 0.406 1577.5 14.125 30001 6.078 30001 70.023 30001
Table 6: Median of CPU time and error for Example 3
$n$ $m$ Alg1 Alg2 Extra
CPU (s) Error CPU (s) Error CPU (s) Error
5 2 0.031 0.01744 0.187 0.01740 0.328 0.01061
5 5 0.148 0.01248 4.234 0.01267 13.632 0.00749
5 10 0.141 0.00920 5.851 0.01658 21.054 0.00608
10 2 0.031 0.02538 0.211 0.02557 1.429 0.01484
10 5 0.078 0.02683 0.992 0.02684 13.828 0.01674
10 10 0.382 0.01663 7.820 0.01658 66.976 0.01062
20 2 0.141 0.15133 0.437 0.01677 6.890 0.01086
20 5 0.679 0.02091 3.031 0.02150 57.523 0.01360
20 10 1.085 0.02167 10.054 0.01919 207.695 0.01195
(a) CPU Time (Alg1-Alg4)
(b) Iteration number (Alg1-Alg4)
(c) CPU Time (Alg1-Alg2 and Extra)
(d) Errors (Alg1-Alg2 and Extra)
Figure 1: First row are comparison among Alg1-Alg4 and the second line are among Alg1-Alg2 and Extra.
5.2 Conclusion
Based on the results of the three types of numerical experiments, we can draw the following observations:
  1. 1.
    From Table 1 and Table 3, we observe that Alg1 and Alg3 perform similarly to Alg2 and Alg4 in terms of CPU time and number of iterations when $m=2$. However, as the number of ellipsoids increases ($m=5,10$), Alg1 and Alg3 significantly outperform Alg2 and Alg4, especially in the case when $m=10$, no matter the dimension $n$. This fact establishes unequivocally the superiority of the circumcenter methods with respect to the older algorithms when the number of ellipsoids increases.
  2. 2.
    From Table 2 and Table 4, the Extragradient method (Extra) demonstrates higher accuracy than both Alg1 and Alg2, which is not surprising, since Extra uses exact projections. Alg2, which includes an inner loop that approximates the exact projection, achieves better accuracy than Alg1. On the other hand, Alg1 achieves significantly faster computation times than Alg2, especially when the number of ellipsoids increases, while maintaining acceptable accuracy. Again, this faster performance confirms the superiority of circumcenter methods.
  3. 3.
    In Example 3, where the operator $F$ is monotone but no longer paramonotone, Alg3 fails to terminate within 30,000 iterations in some cases (see Table 6), consistent with the theoretical requirement of paramonotonicity for its convergence (as indicated in Table 5). In contrast, Alg1, which incorporates the circumcenter-acceleration technique, converges reliably and efficiently, with a reasonable number of iterations. While its accuracy is slightly lower compared to Extra, it is acceptable, and again, it performs much faster than Extra.
  4. 4.
    We also summarize our numerical findings in Fig 1, by means of the so-called performance profiles from [28]. Performance profiles allow one to benchmark different methods on a set of problems with respect to a performance measure (in our case, CPU time and total number of projections). The vertical axis indicates the percentage of problems solved, while the horizontal axis indicates the corresponding factor of the performance index used by the best solver.
In summary, Alg1 and Alg3 demonstrate superior performance over Alg2 and Alg4 in terms of CPU time and iteration count. Although Alg1 is slightly less accurate than Extra, due to the use of approximate projections, it offers a compelling balance between efficiency and accuracy, especially in problems with high-dimensional feasible sets.
6 Concluding remarks
In this work, we propose two distinct circumcenter-type projection methods for solving variational inequality problems involving paramonotone and monotone operators, respectively. Both algorithms enjoy global convergence and ergodic convergence guarantees. Numerical experiments demonstrate that our methods converge significantly faster than existing algorithms, especially when the intersection $\bigcap_{i=1}^{m}C_{i}$ is more complex, while maintaining reasonable accuracy.
Future research may focus on extending these methods to broader classes of operators, such as quasi-monotone, or stochastic settings, and investigating their applications in convex optimization or generalizing them in the equilibrium problem. Additionally, further theoretical analysis on convergence rates, and improving the ergodic convergence to global convergence of Algorithm 2 remain as important issues to be addressed.

### v2

@@ 5.2. Test problems
Finally, we consider monotone operators that are not paramonotone. The operator $F$ is still defined by (5.9) with the block structure (5.10), but now $A_{2}\in\mathds{R}^{n_{2}\times n_{2}}$ is skew-symmetric, i.e.,
(5.11) $A_{2}^{\top}=-A_{2}.$
The operator $F$ is monotone since $\left\langle Ax,x\right\rangle=\left\langle A_{1}x_{1},x_{1}\right\rangle+\left\langle A_{2}x_{2},x_{2}\right\rangle=\left\langle A_{1}x_{1},x_{1}\right\rangle\geq 0$ (as $\left\langle A_{2}x_{2},x_{2}\right\rangle=0$ for skew-symmetric $A_{2}$). However, $F$ is not paramonotone because $\operatorname{rank}(A+A^{\top})<\operatorname{rank}(A)$ when $A_{2}\neq 0$. In this example, we expect that algorithms requiring paramonotonicity (CRM-VIP1 and BI1) may fail to converge in some instances, while BI2, CRM-VIP2, EGM, and Mal-Adap (which only require monotonicity) should converge.
5.3. Discussion on the numerical results
For each example, we present results in tables, in Appendix A, comparing iteration counts and wall-clock time, as well as speedup comparison with the fastest method. Scenarios B and C were combined into a single table for brevity, as the same methods were compared under both scenarios. All the tables report median values over 10 random instances for each configuration of dimension $n$ and number of ellipsoids $m$.
In order to summarize all experiments, in Figures 1, 2, and 3, we present performance profiles [30] for each of the three examples, illustrating the comparative performance of the algorithms across all tested instances. Each figure contains four subplots: the first two correspond to Scenario A (iterations and wall-clock time), while the last two correspond to Scenarios B and C (iterations and wall-clock time). These profiles provide a visual representation of how each algorithm performs relative to the others across different problem instances.
(a) Scenario A: Iterations
(b) Scenario A: Wall-clock time
(c) Scenarios B and C: Iterations
(d) Scenarios B and C: Wall-clock time
Figure 1. Performance profiles for Example 5.1 (gradient operator).
(a) Scenario A: Iterations
(b) Scenario A: Wall-clock time
(c) Scenarios B and C: Iterations
(d) Scenarios B and C: Wall-clock time
Figure 2. Performance profiles for Example 5.2 (paramonotone, non-gradient operator).
(a) Scenario A: Iterations
(b) Scenario A: Wall-clock time
(c) Scenarios B and C: Iterations
(d) Scenarios B and C: Wall-clock time
Figure 3. Performance profiles for Example 5.3 (monotone, non-paramonotone operator).
Based on the results of the three types of numerical experiments reported in Appendix A and summarized in the performance profiles (Figures 13), we draw the following observations:
  1. (1)
    Comparison with Exact Projection Methods (Scenario A): Tables 1, 5, and 9 demonstrate a stark contrast between methods using exact projections (EGM and Mal-Adap) and the proposed approximate methods. CRM-VIP1 and CRM-VIP2 are consistently three to four orders of magnitude faster than EGM and Mal-Adap. For instance, in Table 2 ($n=10,m=2$), CRM-VIP1 is over 10 00010\,000 times faster than EGM. This massive speedup is attributed to the computational cost of projecting onto the intersection of ellipsoids using Dykstra’s algorithm, whereas our methods utilize computationally cheap separating halfspaces. This confirms that for complex feasible sets defined by intersections, approximate projection strategies are computationally superior.
  2. (2)
    Impact of Circumcenter Acceleration (Scenarios A, B, and C): The benefit of the circumcenter acceleration is evident when comparing the proposed methods to their non-accelerated counterparts.
    • CRM-VIP1 vs. BI1: In Scenarios B and C (Tables 311), CRM-VIP1 reduces the iteration count from tens of thousands (often hitting the max-iteration cap) to merely 15–30 iterations. The speedup factors range from roughly $200\times$ to over $3000\times$.
    • CRM-VIP2 vs. BI2: Similarly, CRM-VIP2 exhibits a dramatic improvement over BI2, reducing iteration counts from thousands to single digits (often fewer than 10 iterations). The speedup relative to BI2 frequently exceeds $1000\times$.
    This clearly establishes the superiority of the circumcenter approach, which effectively mitigates the “zig-zagging” behavior often observed in standard separating hyperplane methods.
  3. (3)
    Robustness in Non-Paramonotone Settings (Example 5.3): In Example 5.3, where the operator is monotone but not paramonotone, the theoretical limitations of BI1 become apparent. As shown in Table 11, BI1 fails to converge in some instances (e.g., $n=100,m=5$) or requires an excessive number of iterations. Interestingly, its accelerated version, CRM-VIP1, continues to converge efficiently, suggesting that the circumcenter step may induce a stabilizing effect even outside the strict theoretical scope of paramonotonicity. However, CRM-VIP2, which is theoretically designed for general monotone operators, proves to be the most robust choice, solving all instances in Example 5.3 with the lowest iteration counts (typically under 10).
  4. (4)
    Efficiency vs. Accuracy Trade-off: While EGM and Mal-Adap theoretically provide iterates that lie exactly within the feasible set $C$, the computational cost is prohibitive for large $m$. Both CRM-VIP1 and CRM-VIP2 generate infeasible iterates that approach $C$ asymptotically. However, as shown by the stopping criteria results, the final iterates of the proposed algorithms achieve high accuracy in both feasibility and optimality (solution quality) in a fraction of the time. Between the two proposed methods, CRM-VIP1 is generally faster per iteration (as seen in the speedup tables where CRM-VIP1 is typically $2\times$ to $5\times$ faster than CRM-VIP2), but CRM-VIP2 requires fewer iterations and offers stronger theoretical guarantees for general monotone operators.
  5. (5)
    Global Performance Assessment: The performance profiles in Figures 1, 2, and 3 visually summarize these findings. In all scenarios, the curves for CRM-VIP1 and CRM-VIP2 rise rapidly to 1.0, indicating that they are the most efficient methods on the vast majority of test problems. Specifically, CRM-VIP1 dominates in terms of wall-clock time due to the simplicity of its iteration, while CRM-VIP2 often dominates in terms of iteration count.
In summary, CRM-VIP1 and CRM-VIP2 demonstrate superior performance over BI1, BI2, Mal-Adap, and EGM. We remark that CRM-VIP1 is the fastest method overall, making it the preferred choice for paramonotone problems or gradients of convex functions. On the other hand, CRM-VIP2 offers a compelling alternative for general monotone problems, providing extreme iteration reduction and robustness at a slightly higher cost per iteration than CRM-VIP1.
6. Concluding remarks
In this work, we have proposed two distinct circumcenter-type projection methods for solving variational inequality problems involving paramonotone and monotone operators, respectively. One method is equipped with global convergence guarantees, while the other admits ergodic convergence. The numerical experiments provide compelling evidence of the superiority of the proposed approach. By avoiding computationally expensive exact projections in favor of circumcenter steps onto separating halfspaces, our algorithms achieve a dramatic reduction in computational time—outperforming classical exact projection methods by orders of magnitude and providing robust acceleration over existing approximation schemes, particularly when the intersection $\bigcap_{i=1}^{m}C_{i}$ is complex.
Future research will focus on extending these efficient schemes to broader classes of operators, such as quasi-monotone, or to stochastic settings, as well as investigating their applications in convex optimization and equilibrium problems. Additionally, refining the theoretical analysis of convergence rates and strengthening the ergodic convergence of Algorithm 2 to nonergodic global convergence remain important avenues for future investigation.
Appendix A Tables of numerical results
Table 1. Median of iterations and wall-clock time (s) – Example 5.1.A
$n$ $m$ EGM Mal-Adap BI1 CRM-VIP1 BI2 CRM-VIP2
Iter Time (s) Iter Time (s) Iter Time (s) Iter Time (s) Iter Time (s) Iter Time (s)
5 2 32 $7.21\text{\times}{10}^{-02}$ 50 $6.11\text{\times}{10}^{-02}$ 2065 $2.01\text{\times}{10}^{-03}$ 18 $1.43\text{\times}{10}^{-05}$ 2559 $5.14\text{\times}{10}^{-03}$ 10 $3.06\text{\times}{10}^{-05}$
5 5 37 $1.34\text{\times}{10}^{-01}$ 62 $1.64\text{\times}{10}^{-01}$ 2805 $3.80\text{\times}{10}^{-03}$ 10 $1.58\text{\times}{10}^{-05}$ 5138 $2.39\text{\times}{10}^{-02}$ 4 $2.91\text{\times}{10}^{-05}$
10 2 66 $2.38\text{\times}{10}^{-01}$ 110 $6.07\text{\times}{10}^{-01}$ 3403 $4.14\text{\times}{10}^{-03}$ 20 $2.21\text{\times}{10}^{-05}$ 10538 $5.94\text{\times}{10}^{-02}$ 16 $4.82\text{\times}{10}^{-05}$
10 5 71 $1.61\text{\times}{10}^{-01}$ 69 $1.35\text{\times}{10}^{-01}$ 2094 $3.56\text{\times}{10}^{-03}$ 16 $2.79\text{\times}{10}^{-05}$ 2802 $1.07\text{\times}{10}^{-02}$ 12 $6.28\text{\times}{10}^{-05}$
Table 2. Speedup relative to CRM-VIP1– Example 5.1.A
$n$ $m$ EGM Mal-Adap BI1 BI2 CRM-VIP2
[…]
Table 3. Median of iterations and wall-clock time (s) – Example 5.1.B & C
$n$ $m$ BI1 CRM-VIP1 BI2 CRM-VIP2
Iter Time (s) Iter Time (s) Iter Time (s) Iter Time (s)
50 5 3851 $2.01\text{\times}{10}^{-02}$ 16 $7.21\text{\times}{10}^{-05}$ 9103 $1.57\text{\times}{10}^{-01}$ 10 $1.72\text{\times}{10}^{-04}$
50 8 4906 $3.70\text{\times}{10}^{-02}$ 21 $1.30\text{\times}{10}^{-04}$ 9396 $2.77\text{\times}{10}^{-01}$ 11 $3.02\text{\times}{10}^{-04}$
100 5 3976 $6.40\text{\times}{10}^{-02}$ 28 $2.97\text{\times}{10}^{-04}$ 9340 $5.88\text{\times}{10}^{-01}$ 15 $6.90\text{\times}{10}^{-04}$
100 8 4510 $1.03\text{\times}{10}^{-01}$ 22 $3.36\text{\times}{10}^{-04}$ 8998 $7.53\text{\times}{10}^{-01}$ 12 $8.48\text{\times}{10}^{-04}$
100 20 32712 $1.60\text{\times}{10}^{00}$ 17 $6.04\text{\times}{10}^{-04}$ 1207 $2.49\text{\times}{10}^{-01}$ 5 $1.92\text{\times}{10}^{-03}$
100 30 38040 $2.66\text{\times}{10}^{00}$ 19 $9.66\text{\times}{10}^{-04}$ 1218 $4.51\text{\times}{10}^{-01}$ 9 $3.04\text{\times}{10}^{-03}$
100 50 41405 $4.68\text{\times}{10}^{00}$ 16 $1.20\text{\times}{10}^{-03}$ 1600 $8.26\text{\times}{10}^{-01}$ 7 $5.10\text{\times}{10}^{-03}$
200 20 31868 $6.86\text{\times}{10}^{00}$ 24 $2.41\text{\times}{10}^{-03}$ 1323 $1.28\text{\times}{10}^{00}$ 13 $1.03\text{\times}{10}^{-02}$
200 30 29778 $9.80\text{\times}{10}^{00}$ 21 $2.80\text{\times}{10}^{-03}$ 1266 $2.21\text{\times}{10}^{00}$ 11 $1.43\text{\times}{10}^{-02}$
200 50 23601 $1.18\text{\times}{10}^{01}$ 20 $3.95\text{\times}{10}^{-03}$ 984 $2.22\text{\times}{10}^{00}$ 6 $2.18\text{\times}{10}^{-02}$
500 20 17461 $1.40\text{\times}{10}^{01}$ 22 $1.08\text{\times}{10}^{-02}$ 851 $4.22\text{\times}{10}^{00}$ 12 $4.43\text{\times}{10}^{-02}$
500 30 14873 $1.72\text{\times}{10}^{01}$ 17 $1.05\text{\times}{10}^{-02}$ 701 $5.13\text{\times}{10}^{00}$ 7 $5.71\text{\times}{10}^{-02}$
500 50 978 $2.51\text{\times}{10}^{00}$ 32 $6.15\text{\times}{10}^{-02}$ 549 $5.68\text{\times}{10}^{00}$ 6 $1.58\text{\times}{10}^{-01}$
Table 4. Speedup relative to CRM-VIP1– Example 5.1.B & C
$n$ $m$ BI1 BI2 CRM-VIP2
[…]
Table 5. Median of iterations and wall-clock time (s) – Example 5.2.A
$n$ $m$ EGM Mal-Adap BI1 CRM-VIP1 BI2 CRM-VIP2
Iter Time (s) Iter Time (s) Iter Time (s) Iter Time (s) Iter Time (s) Iter Time (s)
5 2 45 $7.26\text{\times}{10}^{-02}$ 38 $26\text{\times}{10}^{-01} 11$54 $1.04\text{\times}{10}^{-03}$ 19 $1.45\text{\times}{10}^{-05}$ 1759 $3.20\text{\times}{10}^{-03}$ 5 $3.09\text{\times}{10}^{-05}$
5 5 48 $1.87\text{\times}{10}^{-01}$ 62 $3.16\text{\times}{10}^{-01}$ 2566 $3.22\text{\times}{10}^{-03}$ 15 $2.27\text{\times}{10}^{-05}$ 10158 $9.56\text{\times}{10}^{-02}$ 6 $4.54\text{\times}{10}^{-05}$
10 2 34 $1.75\text{\times}{10}^{-01}$ 60 $1.54\text{\times}{10}^{-01}$ 4112 $4.47\text{\times}{10}^{-03}$ 18 $1.85\text{\times}{10}^{-05}$ 6378 $1.39\text{\times}{10}^{-02}$ 5 $3.17\text{\times}{10}^{-05}$
10 5 40 $2.04\text{\times}{10}^{-01}$ 67 $3.73\text{\times}{10}^{-01}$ 2440 $4.93\text{\times}{10}^{-03}$ 14 $3.06\text{\times}{10}^{-05}$ 10108 $7.54\text{\times}{10}^{-02}$ 6 $5.74\text{\times}{10}^{-05}$
Table 6. Speedup relative to CRM-VIP1– Example 5.2.A
$n$ $m$ EGM Mal-Adap BI1 BI2 CRM-VIP2
[…]
Table 7. Median of iterations and wall-clock time (s) – Example 5.2.B & C
$n$ $m$ BI1 CRM-VIP1 BI2 CRM-VIP2
Iter Time (s) Iter Time (s) Iter Time (s) Iter Time (s)
50 5 5131 $3.90\text{\times}{10}^{-02}$ 19 $17\text{\times}{10}^{-04} 11$904 $3.01\text{\times}{10}^{-01}$ 11 $2.38\text{\times}{10}^{-04}$
50 8 4971 $5.07\text{\times}{10}^{-02}$ 26 $2.21\text{\times}{10}^{-04}$ 9198 $3.84\text{\times}{10}^{-01}$ 12 $4.30\text{\times}{10}^{-04}$
100 5 3738 $7.92\text{\times}{10}^{-02}$ 18 $2.41\text{\times}{10}^{-04}$ 8226 $6.05\text{\times}{10}^{-01}$ 8 $5.63\text{\times}{10}^{-04}$
100 8 3331 $8.52\text{\times}{10}^{-02}$ 13 $2.33\text{\times}{10}^{-04}$ 9212 $8.77\text{\times}{10}^{-01}$ 4 $4.87\text{\times}{10}^{-04}$
100 20 2266 $1.34\text{\times}{10}^{-01}$ 16 $8.64\text{\times}{10}^{-04}$ 1108 $3.15\text{\times}{10}^{-01}$ 7 $3.89\text{\times}{10}^{-03}$
100 30 38772 $3.30\text{\times}{10}^{00}$ 17 $1.18\text{\times}{10}^{-03}$ 1209 $4.87\text{\times}{10}^{-01}$ 7 $5.05\text{\times}{10}^{-03}$
100 50 1054 $1.49\text{\times}{10}^{-01}$ 15 $1.49\text{\times}{10}^{-03}$ 1291 $5.43\text{\times}{10}^{-01}$ 6 $7.00\text{\times}{10}^{-03}$
200 20 24715 $6.41\text{\times}{10}^{00}$ 16 $3.76\text{\times}{10}^{-03}$ 1075 $1.24\text{\times}{10}^{00}$ 7 $1.19\text{\times}{10}^{-02}$
200 30 25707 $9.24\text{\times}{10}^{00}$ 21 $5.72\text{\times}{10}^{-03}$ 1113 $2.41\text{\times}{10}^{00}$ 8 $2.57\text{\times}{10}^{-02}$
200 50 822 $4.05\text{\times}{10}^{-01}$ 22 $7.90\text{\times}{10}^{-03}$ 867 $1.44\text{\times}{10}^{00}$ 5 $3.76\text{\times}{10}^{-02}$
500 20 18475 $2.30\text{\times}{10}^{01}$ 19 $1.58\text{\times}{10}^{-02}$ 813 $6.68\text{\times}{10}^{00}$ 9 $6.81\text{\times}{10}^{-02}$
500 30 13732 $2.58\text{\times}{10}^{01}$ 12 $1.56\text{\times}{10}^{-02}$ 738 $7.18\text{\times}{10}^{00}$ 3 $5.19\text{\times}{10}^{-02}$
500 50 16452 $5.33\text{\times}{10}^{01}$ 33 $6.73\text{\times}{10}^{-02}$ 799 $1.45\text{\times}{10}^{01}$ 6 $3.29\text{\times}{10}^{-01}$
Table 8. Speedup relative to CRM-VIP1– Example 5.2.B & C
$n$ $m$ BI1 BI2 CRM-VIP2
[…]
Table 9. Median of iterations and wall-clock time (s) – Example 5.3.A
$n$ $m$ EGM Mal-Adap BI1 CRM-VIP1 BI2 CRM-VIP2
Iter Time (s) Iter Time (s) Iter Time (s) Iter Time (s) Iter Time (s) Iter Time (s)
5 2 58 $1.93\text{\times}{10}^{-01}$ 34 $2.51\text{\times}{10}^{-01}$ 2569 $3.02\text{\times}{10}^{-03}$ 17 $1.89\text{\times}{10}^{-05}$ 9549 $4.52\text{\times}{10}^{-02}$ 11 $3.57\text{\times}{10}^{-05}$
5 5 53 $1.41\text{\times}{10}^{-01}$ 45 $2.92\text{\times}{10}^{-01}$ 4256 $7.02\text{\times}{10}^{-03}$ 11 $2.10\text{\times}{10}^{-05}$ 14770 $8.83\text{\times}{10}^{-02}$ 6 $5.18\text{\times}{10}^{-05}$
10 2 29 $2.60\text{\times}{10}^{-01}$ 56 $5.24\text{\times}{10}^{-01}$ 6978 $9.27\text{\times}{10}^{-03}$ 12 $1.46\text{\times}{10}^{-05}$ 15400 $8.00\text{\times}{10}^{-02}$ 4 $2.94\text{\times}{10}^{-05}$
10 5 26 $3.33\text{\times}{10}^{-01}$ 47 $6.35\text{\times}{10}^{-01}$ 4815 $9.49\text{\times}{10}^{-03}$ 12 $2.70\text{\times}{10}^{-05}$ 12973 $1.06\text{\times}{10}^{-01}$ 4 $5.14\text{\times}{10}^{-05}$
Table 10. Speedup relative to CRM-VIP1– Example 5.3.A
$n$ $m$ EGM Mal-Adap BI1 BI2 CRM-VIP2
[…]
Table 11. Median of iterations and wall-clock time (s) – Example 5.3.B & C
$n$ $m$ BI1 CRM-VIP1 BI2 CRM-VIP2
Iter Time (s) Iter Time (s) Iter Time (s) Iter Time (s)
50 5 5647 $4.29\text{\times}{10}^{-02}$ 16 $9.55\text{\times}{10}^{-05}$ 13759 $4.13\text{\times}{10}^{-01}$ 10 $2.13\text{\times}{10}^{-04}$
50 8 6027 $6.18\text{\times}{10}^{-02}$ 20 $1.69\text{\times}{10}^{-04}$ 13053 $6.40\text{\times}{10}^{-01}$ 8 $3.45\text{\times}{10}^{-04}$
100 5 10 $1.34\text{\times}{10}^{-04}$ 10839 $1.14\text{\times}{10}^{00}$ 4 $2.52\text{\times}{10}^{-04}$
100 8 5298 $1.68\text{\times}{10}^{-01}$ 20 $3.61\text{\times}{10}^{-04}$ 10900 $1.36\text{\times}{10}^{00}$ 9 $8.28\text{\times}{10}^{-04}$
100 20 32921 $2.02\text{\times}{10}^{00}$ 16 $8.94\text{\times}{10}^{-04}$ 1346 $3.84\text{\times}{10}^{-01}$ 8 $3.57\text{\times}{10}^{-03}$
100 30 38636 $3.60\text{\times}{10}^{00}$ 18 $1.27\text{\times}{10}^{-03}$ 1573 $6.56\text{\times}{10}^{-01}$ 6 $4.95\text{\times}{10}^{-03}$
100 50 40032 $4.64\text{\times}{10}^{00}$ 16 $9.25\text{\times}{10}^{-04}$ 1678 $1.07\text{\times}{10}^{00}$ 8 $4.56\text{\times}{10}^{-03}$
200 20 33695 $7.84\text{\times}{10}^{00}$ 22 $2.72\text{\times}{10}^{-03}$ 1394 $2.08\text{\times}{10}^{00}$ 9 $1.12\text{\times}{10}^{-02}$
200 30 32940 $1.07\text{\times}{10}^{01}$ 15 $2.23\text{\times}{10}^{-03}$ 1228 $3.21\text{\times}{10}^{00}$ 7 $1.18\text{\times}{10}^{-02}$
200 50 23959 $1.25\text{\times}{10}^{01}$ 15 $2.95\text{\times}{10}^{-03}$ 1198 $3.13\text{\times}{10}^{00}$ 4 $1.24\text{\times}{10}^{-02}$
500 20 18383 $1.52\text{\times}{10}^{01}$ 23 $1.60\text{\times}{10}^{-02}$ 864 $6.97\text{\times}{10}^{00}$ 7 $6.30\text{\times}{10}^{-02}$
500 30 19819 $2.78\text{\times}{10}^{01}$ 17 $1.79\text{\times}{10}^{-02}$ 947 $1.13\text{\times}{10}^{01}$ 7 $6.88\text{\times}{10}^{-02}$
500 50 15070 $3.92\text{\times}{10}^{01}$ 9 $1.67\text{\times}{10}^{-02}$ 876 $1.74\text{\times}{10}^{01}$ 3 $5.15\text{\times}{10}^{-02}$
Table 12. Speedup relative to CRM-VIP1– Example 5.3.B & C
$n$ $m$ BI1 BI2 CRM-VIP2
[…]